# SKEL-CF: Coarse-to-Fine Biomechanical Skeleton and Surface Mesh Recovery

### v3

@@ 4.4 Discussion
To analyze the contribution of each component, we conduct an ablation study on the MOYO [moyo] and COCO [coco] datasets, representing 3D and 2D evaluation metrics, respectively. We progressively activate three key components: the camera model, the coarse-to-fine strategy, and the refinement stage, then report the corresponding performance under both 3D and 2D metrics. Incorporating the camera model notably improves the 3D reconstruction accuracy. The coarse-to-fine strategy further enhances the 2D metrics. Finally, the refinement stage brings consistent improvements across both 3D and 2D evaluations.
Iterative mesh refinement process.
We further visualize the iterative mesh refinement process of SKEL-CF in Figure 5. As shown, our model begins with a coarse estimation of the motion and progressively refines it through successive iterations, leading to a more accurate reconstruction of complex motions. This visualization clearly demonstrates the effectiveness of the proposed coarse-to-fine refinement strategy. Additional video results illustrating the refinement process are provided in the supplementary material.
@@
PyMAF [pymaf2021] 0.68 0.86
CLIFF [li2022cliff] 0.64 0.88
PARE [kocabas2021pare] 0.72 0.91
PyMAF-X [pymafx2023] 0.79 0.93
HMR2.0a [goel20234dhuman] 0.79 0.94
CameraHMR [patel2024camerahmr] 0.84 0.94
HSMR [xia2025hsmr] 0.85 0.96
@@
Table 3: 2D Keypoint Performance on COCO [coco]. For completeness, we also report PCK scores of the projected keypoints at thresholds of 0.05 and 0.1, showing that SKEL-CF achieves comparable 2D keypoint quality while maintaining significant better biomechanical mesh quality.
Components MOYO [moyo] COCO [coco]
@@
Table 4: Ablation study on MOYO [moyo] and COCO [coco]. Cam denotes the camera model, C2F represents the coarse-to-fine strategy, and Refine indicates the progressive refinement stage.
@@ 5 Conclusion
In this work, we introduced SKEL-CF, a comprehensive biomechanical skeleton and surface mesh recovery pipeline built upon the SKEL representation. By transforming the high-quality SMPL dataset into a SKEL-based format, termed 4DHuman-SKEL, we established a strong foundation for accurate and generalizable training. Our integration of an explicit camera model effectively mitigates the limitations of weak-perspective assumptions seen in previous works such as HSMR, enabling robust pose estimation under diverse viewpoints. Furthermore, the proposed coarse-to-fine refinement strategy enables stable and high-fidelity reconstruction results. Overall, SKEL-CF sets a new baseline for biomechanical skeleton and surface recovery, providing a practical and extensible framework for future research in human modeling and motion understanding.
@@
Supplementary Material
S1 Organization
This supplementary document is organized as follows:
  • This supplementary material includes additional video demonstrations showcasing both our SKEL-CF visualizations and comparative results with HSMR [xia2025hsmr] and CameraHMR [patel2024camerahmr] (located in the videos-demo folder). It also provides extended coarse-to-fine visualizations (in the coarse-to-fine-gif folder) and the supplementary PDF (SKEL-CF-sup.pdf).
  • Additional qualitative results comparing our method with the SKEL-based model HSMR [xia2025hsmr] are presented in Fig. S2.
  • Additional qualitative comparisons with the SMPL [loper2015smpl]-based approach CameraHMR [patel2024camerahmr] are provided in Fig S1.
  • Details on the MOYO-HARD subset are in Section S6, see line 380 of the main text.
  • More visual results illustrating our coarse-to-fine refinement strategy are shown in Fig. S3, as a supplement to Fig. 5 in the main text (line 468).
S2 Visual comparison with SMPLs
To more comprehensively evaluate our model, we present in Fig. S1 additional visual comparisons extending the results shown in Fig. 4 of the main paper. These examples further compare our method with the state-of-the-art SMPL [loper2015smpl]-based approach, CameraHMR [patel2024camerahmr]. It indicates that our method can reconstruct more natural pose with accurate anatomical joints.
S3 Visual comparison with SKELs
To further compare SKEL [SKEL]-based approaches, Fig. S2 shows additional qualitative comparisons with the state-of-the-art HSMR [xia2025hsmr] pipeline. Unlike HSMR, our method incorporates explicit camera modeling, enabling more reliable and accurate pose reconstruction under varying viewpoints.
S4 More results of Coarse-to-Fine
Figure S3 presents additional visual examples from LSP-ext [lspnet], highlighting how the refinement process progressively improves pose and shape estimation.
S5 Per-Layer Attention Analysis
S6 Details of MOYO-HARD dataset
The MOYO [moyo] test set consists of 29 yoga sequences, each recorded simultaneously by eight cameras capturing the same pose from different viewpoints. Each sequence progresses from simple poses to complex ones and then returns to simpler poses, as shown in Figure S5. To specifically analyze performance under challenging poses, we focus on the middle portion of each sequence, where the motions are most complex. Accordingly, we construct the MOYO-HARD subset by selecting the middle 50% of frames from videos captured by Camera #1 (front view), resulting in a total of 9,725 images.
Table S1: COCO results under different camera extrinsic factorizations ($s$, $t_{x}$, $t_{y}$). (A) Fix final SKEL [SKEL] parameters and $s$, vary layer-wise $(t_{x},t_{y})$. (B) Fix final SKEL parameters and $(t_{x},t_{y})$, vary layer-wise $s$.
@@
Figure S1: Additional comparisons with SMPL [loper2015smpl]-based models. Compared with the state-of-the-art SMPL-based method CameraHMR [patel2024camerahmr], our approach produces more natural and anatomically consistent joint predictions.
Figure S2: Additional comparisons with SKEL [SKEL]-based models. Compared with the state-of-the-art HSMR [xia2025hsmr] method, our approach predicts human poses that are more realistic and accurate.
Figure S3: Additional coarse-to-fine refinement results on LSP-ext [lspnet]. These examples illustrate how our coarse-to-fine strategy progressively improves pose and shape estimation, yielding more accurate and stable reconstructions across diverse human poses.

### v4

@@ 4.4 Discussion
To analyze the contribution of each component, we conduct an ablation study on MOYO [33] and COCO [22], representing 3D and 2D evaluation metrics, respectively. We start from a baseline that removes camera modeling, coarse-to-fine (C2F), and iterative refinement, which is architecturally comparable to HSMR [39]. In this controlled setting, the only difference from HSMR is the training data. The large performance gap on MOYO (e.g., PA-MPJPE: 53.7 vs. 79.6) highlights the effectiveness of our higher-quality training annotations. Building on this baseline, we progressively activate three key components: an explicit camera model, the coarse-to-fine strategy, and the iterative refinement stage, and report the corresponding performance under both 3D and 2D metrics. Incorporating the camera model notably improves 3D reconstruction accuracy. The coarse-to-fine strategy further enhances the 2D metrics. Finally, the iterative refinement stage brings consistent improvements across both 3D and 2D evaluations.
Iterative mesh refinement.
We further visualize the iterative mesh refinement process of SKEL-CF in Figure 5. This visualization clearly demonstrates the effectiveness of the proposed coarse-to-fine refinement strategy. Additional visual results illustrating the refinement process are provided in Appendix S2.
@@
PyMAF [43] 0.68 0.86
CLIFF [21] 0.64 0.88
PARE [16] 0.72 0.91
PyMAF-X [42] 0.79 0.93
HMR2.0a [6] 0.79 0.94
CameraHMR [27] 0.84 0.94
HSMR [39] 0.85 0.96
@@
Table 3: 2D Keypoint Performance on COCO [22]. For completeness, we also report PCK scores of the projected keypoints at thresholds of 0.05 and 0.1, showing that SKEL-CF achieves comparable 2D keypoint quality while maintaining significant better biomechanical mesh quality.
Components MOYO [33] COCO [22]
@@
Table 4: Ablation study on MOYO [33] and COCO [22]. Cam denotes the camera model, C2F represents the coarse-to-fine strategy, and Refine indicates the iterative refinement stage.
@@ 5 Conclusion
In this work, we introduced SKEL-CF, a biomechanical skeleton and surface mesh recovery pipeline built upon the SKEL representation. By transforming the high-quality SMPL dataset into a SKEL-based format, termed 4DHuman-SKEL, we established a strong foundation for accurate and generalizable training. Our integration of an explicit camera model effectively mitigates the limitations of weak-perspective assumptions seen in previous works such as HSMR, enabling robust pose estimation under diverse viewpoints. Furthermore, the proposed coarse-to-fine refinement strategy enables stable and high-fidelity reconstruction results. Overall, SKEL-CF sets a new baseline for biomechanical skeleton and surface recovery, providing a practical and extensible framework for future research in human modeling and motion understanding.
Appendix
S1 Organization
This appendix document is organized as follows:
  • Additional video demonstrations showcasing both our SKEL-CF visualizations and comparative results with HSMR [39] and CameraHMR [27] are provided in the project page: https://pokerman8.github.io/SKEL-CF/
  • Details on the iterative refinement are in Section S2, more visual results illustrating our coarse-to-fine refinement strategy are shown in Fig. S4, as a supplement to Fig. 5.
  • Implementation details are in Section S3.
  • Additional qualitative results comparing our method with the SKEL-based model HSMR [39] are presented in Fig. S3.
  • Additional qualitative comparisons with the SMPL [23]-based approach CameraHMR [27] are provided in Fig S2.
  • Details on the MOYO-HARD subset are in Section S4.
S2 Details of iterative refinement
Figure S4 presents additional visual examples from LSP-ext [10], highlighting how the refinement process progressively improves pose and shape estimation. We further validate this design in Table S1: the proposed sparse refinement supervision consistently improves performance on MOYO and MOYO-HARD and achieves accuracy comparable to full per-layer prediction, while being substantially more efficient. In addition, repeating a single decoder layer degrades performance, especially on MOYO-HARD, highlighting the importance of progressive refinement across layers.
Method Loss Decoder Layer 3DPW MOYO MOYO-HARD
@@
Sparse super. (Ours) $\mathcal{L}_{refine}$ 6 $\times$ 1 61.5 38.7 73.5 85.0 51.4 91.9 90.0 61.5 102.5
Full predict. $\mathcal{L}_{skel}$ 6 $\times$ 1 61.3 38.6 73.3 86.4 53.2 93.9 92.1 65.2 105.6
One layer $\mathcal{L}_{refine}$ 1 $\times$ 6 62.5 40.1 74.8 86.8 54.4 94.2 96.8 70.1 111.5
Table S1: Further study of iterative refinement. Sparse super. applies refinement supervision across decoder layers, Full predict. predicts full SKEL parameters at every layer, and One layer repeats a single layer for multiple refinement steps. “Decoder Layer $\times$ Refine Time” summarizes the refinement configuration.
Table S2: COCO results under different camera extrinsic factorizations ($s$, $t_{x}$, $t_{y}$). (A) Fix final SKEL [15] parameters and $s$, vary layer-wise $(t_{x},t_{y})$. (B) Fix final SKEL parameters and $(t_{x},t_{y})$, vary layer-wise $s$.
@@
S3 Implementation details
Following HSMR [39], we adopt ViTPose-H [41], pretrained on COCO [22], as the encoder backbone. ViTPose-H consists of 32 transformer layers, each with 16 attention heads and a hidden dimension of 1280. The decoder is a standard transformer decoder with $L=6$ layers, consistent with the design in HSMR [39]. We train SKEL-CF using the AdamW [24] optimizer with $\beta_{1}$, $\beta_{2}$ setting to 0.9, 0.999, weight decay setting to $1\times 10^{-4}$, a batch size of 64 and a learning rate of $1\times 10^{-5}$, preceded by a one-epoch warm-up. Training is conducted for 30 epochs on 8 NVIDIA A100 GPUs, taking approximately 120 hours in total. The hyper-parameters are set as $\lambda_{kp}=0.05$, $\lambda_{\beta}=0.0005$, $\lambda_{\theta}=0.001$, and $\lambda_{ref}=0.1$. Notably, our training schedule is significantly shorter than that of HSMR [39], which trains for 100 epochs.
S4 Details of MOYO-HARD dataset
The MOYO [33] test set consists of 29 yoga sequences, each recorded simultaneously by eight cameras capturing the same pose from different viewpoints. Each sequence progresses from simple poses to complex ones and then returns to simpler poses, as shown in Figure S1. To specifically analyze performance under challenging poses, we focus on the middle portion of each sequence, where the motions are most complex. Accordingly, we construct the MOYO-HARD subset by selecting the middle 50% of frames from videos captured by Camera #1 (front view), resulting in a total of 9,725 images.
S5 Visual comparison with SMPLs
To more comprehensively evaluate our model, we present in Fig. S2 additional visual comparisons extending the results shown in Fig. 4 of the main paper. These examples further compare our method with the state-of-the-art SMPL [23]-based approach, CameraHMR [27]. It indicates that our method can reconstruct more natural pose with accurate anatomical joints.
S6 Visual comparison with SKELs
To further compare SKEL [15]-based approaches, Fig. S3 shows additional qualitative comparisons with the state-of-the-art HSMR [39] pipeline. Unlike HSMR, our method incorporates explicit camera modeling, enabling more reliable and accurate pose reconstruction under varying viewpoints.
S7 Per-Layer Attention Analysis
Figure S2: Additional comparisons with SMPL [23]-based models. Compared with the state-of-the-art SMPL-based method CameraHMR [27], our approach produces more natural and anatomically consistent joint predictions.
Figure S3: Additional comparisons with SKEL [15]-based models. Compared with the state-of-the-art HSMR [39] method, our approach predicts human poses that are more realistic and accurate.
Figure S4: Additional iterative refinement results on LSP-ext [10]. These examples illustrate how our iterative refinement strategy progressively improves pose and shape estimation, yielding more accurate and stable reconstructions across diverse human poses.